OpenStudy (anonymous):

Find the value of 3 over the square root of 3..

OpenStudy (anonymous):

thats 27 3*3*3 = 27

myininaya (myininaya):

\[\frac{3}{\sqrt{3}} \cdot \frac{\sqrt{3}}{\sqrt{3}} \]

OpenStudy (anonymous):

^that's called 'rationalizing' :)

OpenStudy (anonymous):

There are 2 ways to go about this question, depending on what the question really wants. You can punch everything into the calculator to give you an approximate value, 3/ sqrt(3) = 1.73 approx. Or the question actually requires you to rationalize the expression. Given 3/ sqrt(3), you multiply sqrt(3) to both the numerator and denominator, giving just sqrt(3).

OpenStudy (anonymous):

A. 3 B. 3 times the square root of 3. C. the square root of 3. D. the square root of 3 over 3. ???

OpenStudy (anonymous):

c) but do you understand how you got the answer?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

Find the value of 10 times the square root of 2 over the square root of 5. Choose one answer. A. 2 B. 2 times the square root of 5. C. 2 the square root of 10. D. 2 over 5 times the square root of 10.

OpenStudy (anonymous):

Just like before you need to rationalize it, do you remember how to rationalize?

OpenStudy (anonymous):

the answer is B RIGHT

OpenStudy (anonymous):

now you are just guessing, and that's not good. If you have a square root in the denominator, do you know how to get it into the numerator: hint: same as the last question you asked.

OpenStudy (anonymous):

rAtionalize

OpenStudy (anonymous):

yup, so how would you rationalise? \[\frac{10\sqrt{2}}{\sqrt{5}} \]

OpenStudy (anonymous):

hint: you need to multiply it by something. Same way we did in the last question

OpenStudy (anonymous):

still on this question?

OpenStudy (anonymous):

yes 2 diveded by 10 than 5 times 5?

OpenStudy (anonymous):

@dahmanman you gave him the wrong answer to the first question

OpenStudy (anonymous):

hats the answer than

OpenStudy (anonymous):

@suzukirider1400 yes! 10 square root 2 divided by square root 5 times square root 5 divided by the square root five, or: \[\frac{10\sqrt{2}}{\sqrt{5}} \times \frac{\sqrt{5}}{\sqrt{5}}\] so now what should we do?

OpenStudy (anonymous):

let's start in the denominator. do you know what: \[\sqrt{5} \times \sqrt{5} = ?\]